The owner of a health foods store sells raisins for $3.20 per pound and nuts for $2.40 per pound. He decides to mix the raisins and the nuts and sell the mixture for $2.72 per pound. To get 50 pounds of the mixture how many pounds of raisins and of nuts should he use?
Answer by jorel1380(3719)

50x2.72=136
p+q=50
3.2p+2.4q=136
2.4p+2.4q=120
.8p=16
p=16/.8=20
q=30
The owner needs 20 lbs. of $3.20 mix and 30 lbs. of $2.40 mix..